A rectangle has a perimeter of 46 m. The length of the rectangle 3 times longer than its width. What is the width of the rectangle.
Answer by Alan3354(69443)

A rectangle has a perimeter of 46 m. The length of the rectangle 3 times longer than its width. What is the width of the rectangle.
------------
1 time longer = 2 times as long
2 times longer = 3 times as long
3 times longer = 4 times as long
---
P = 2L + 2W = 46
L + W = 23
L = 4W
4W + W = 23
4W = 23
W = 4.6 meters